What is the equation of the line that passes through the points (8,-4) and (6,4)?
Answer by Shana-D77(132)

When given two points and asked to write the equation, always find slope (m) first.


(8,-4) and (6,4)
slope (m) = (y2 - y1) / (x2 - x1)


(8, -4) is (x1, y1)
(6, 4) is (x2, y2)


m = (4 - -4) / (6 - 8)
m = 8/-2
m = -4


When given 2 points, like you are, after finding the slope you get to pick your favorite point. My favorite from above is (6, 4) because negative numbers stink! So x1 = 6 and y1 = 4.


Now we use a very special formula, the point-slope formula:
y - y1 = m(x - x1)
In this formula, y and x always stay as y and x. m is the slope you just found. Now we'll plug in m = -4 through point (6, 4):


y - 4= -4(x - 6)
y - 4 = -4x + 24 (distributed the -4)
y = -4x + 28 (added 4 to both sides)